# Measuring deviations from spherical symmetry

### v1

@@ 3.2 Some statistical applications
By Proposition 3.1, Theorem 3.1 and 3.2, the statistic $\hat{\mathcal{M}}_{n}^{2}$ is a consistent estimator of the measure $\mathcal{M}^{2}$, which defines the deviation from sphericity. To use these results for uncertainty quantification, one requires estimates of the variances $s_{n}^{2}$ and $\sigma^{2}$ defined in (3.10) and (3.12), respectively.
@@
which is obtained by replacing integral of the squared density in (3.10) by its corresponding estimate (2.8). We will use this estimate to define a test for exact sphericity (see Remark 3.2 below).
The estimation of $\sigma^{2}$ is more difficult and we propose a jackknife approach for this purpose (see Chapter 5 of Lee,, 2003)). To be precise, let $\hat{\mathcal{M}}^{2}_{n-1}(-i)$ denote the estimator (2.12) of the minimum distance ${\cal M}^{2}$ based on the observations $Y_{1},\ldots,Y_{i-1},Y_{i+1},\ldots,Y_{n}$, and define the pseudovalues
@@
denotes the jackknife estimate of the mean of $\mathcal{M}_{n}^{2}$.
@@
where $\Delta>0$ is a prespecified threshold. Note that this perspective of hypothesis testing aligns with the view expressed by Berger and Delampady, (1987), who argue that it is rare, and perhaps impossible, to have a null hypothesis that can be exactly modeled by a parameter being precisely 0. Similarly, Tukey, (1991), in the context of multiple comparisons of means, emphasizes that “All we know about the world teaches us that the effects of A and B are always different — in some decimal place — for any A and B. Thus, asking ‘Are the effects different?’ is foolish.” We call hypotheses the form (3.24) and (3.25) relevant hypotheses in the following discussion. Note also that hypotheses of the form (3.25) have found considerable attention in equivalence testing in the field of biostatistics, which explains the notations $H_{0}^{\rm eq}$ and $H_{1}^{\rm eq}$ in (3.25) (see Wellek,, 2010) and that rejection of the null hypothesis in (3.25) allows to decide for approximate sphericity at a controlled type I error.
@@
    $\lim_{n\rightarrow\infty}\mathbb{P}\Big(\hat{\mathcal{M}}_{n}^{2}>\Delta+u_{1-\alpha}{\hat{s}_{n}\over\sqrt{n}}\Big)=\begin{cases}1,&\text{if }{\mathcal{M}}^{2}>\Delta,\\ \alpha,&\text{if }{\mathcal{M}}^{2}=\Delta,\\ 0,&\text{if }{\mathcal{M}}^{2}<\Delta.\end{cases}$ (3.28)
@@
    $\lim_{n\rightarrow\infty}\mathbb{P}\Big(\hat{\mathcal{M}}_{n}^{2}\leq\Delta+u_{\alpha}{\hat{s}_{n}\over\sqrt{n}}\Big)=\begin{cases}1,&\text{if }{\mathcal{M}}^{2}<\Delta,\\ \alpha,&\text{if }{\mathcal{M}}^{2}=\Delta,\\ 0,&\text{if }{\mathcal{M}}^{2}>\Delta.\end{cases}$ (3.29)
@@
    An important question in testing hypotheses of the form (3.24) and (3.25) is the choice of the threshold $\Delta$, a similar argument can be given for the hypotheses (3.24). Note that this can be done in a data adaptive way. To be precise we concentrate on the hypotheses (3.25), and note these hypotheses are nested, that is if $H_{0}^{\rm eq}$ is satisfied for $\Delta_{1}$, then it is also satisfied for all $\Delta_{2}\leq\Delta_{1}$. The decision rules for different $\Delta$’s are nested in the same (if $H_{0}^{\rm eq}$ is rejected for $\Delta_{2}$, it also rejected for all $\Delta_{1}\geq\Delta_{2}$). By the sequential rejection principle, we may simultaneously test the hypotheses in (3.25) for different $\Delta\geq 0$ starting at $\Delta=0$ and increasing $\Delta$ to find the minimum value of $\Delta$, say
    $\hat{\Delta}_{\alpha}:=\min\big\{\Delta\geq 0\,|\,\hat{\cal M}^{2}_{n}\leq\Delta+u_{\alpha}\hat{s}_{n}/\sqrt{n}\big\}$ (3.30)
@@
    it is worthwhile to mention that the theory developed so far also provides a test for these hypotheses. More precisely, we propose to reject the null hypothesis in (3.31), whenever
@@
    where $\hat{s}_{n}$ is defined in (3.19). It then follows from Theorem 3.1 and 3.2 that this decision rule defines a consistent and asymptotic level $\alpha$ test.
@@ 4 Pivotal inference
The statistical methodology developed in Section 3.2 requires estimation of the variance $\hat{s}_{n}^{2}$. In this section, we develop a pivotal confidence interval for the deviation $\mathcal{M}^{2}$ from sphericity and pivotal tests for the hypotheses (3.24) and (3.25). For this purpose, we prove a weak convergence result for a sequential version of the estimator $\hat{\mathcal{M}}_{n}^{2}$.
@@
whose asymptotic properties are investigated in following theorem. Throughout this section the symbol $\Rightarrow$ denotes weak convergence in the space $\ell^{\infty}([0,1])$ of bounded functions on the interval $[0,1]$.
@@
We will now use the result (4.3) to develop pivotal inference for the measure $\mathcal{M}^{2}$. To be precise, we define the statistic
@@
Similarly, we propose as a test for the hypotheses (3.24) and (3.25) to reject the null hypothesis, whenever
@@ 5 Finite sample properties
In this section we illustrate the finite sample properties of the developed methodology by means of a small simulation study. All results presented here are based on $1000$ simulation runs.
According to Theorem 3.2, the term $\kappa^{(p-1)/2}/(nh)$ controls the variance for the non-leading term in the Gaussian approximation, while $\sqrt{n}/\kappa^{2}+\sqrt{n}h^{4}$ accounts for the bias term with the bias reduction procedure discussed in Remark 3.2. In addition, for the sequential convergence, by Theorem 4.1, the inflated rate $(\log n)^{2}\kappa^{(p-1)/2}/(nh)$ controls the variance of the process of the non-leading term. In order to approximate Gaussianity in finite samples better, we recommend imposing a stronger control over the non-leading term, that is, $n^{3/2}\kappa^{(p-1)/2}/(nh)\to 0$. Therefore, we consider bandwidths $(h,\kappa)$ satisfying
@@
In the following discussion we consider two models.
@@
The deviation $\mathcal{M}^{2}$ was calculated numerically as $\mathcal{M}^{2}\approx 0.95$ , using a using a larger sample size $2500$ and averaged over $200000$ times of repeated simulation. For the selection of the smoothing parameters via (5.2), we use
@@ 5.1 Confidence intervals for the deviation from a spherical distribution
For Model 1, the empirical coverage rates of both methods are close to their nominal levels for sample sizes larger than $500$. The Jackknife confidence interval achieves more accurate simulated coverage rates when the sample size is small, while the pivotal confidence intervals are more conservative for smaller sample sizes. A potential explanation of this observation is that in this case the widths of the pivotal method are much larger than those of the Jackknife method.
@@ 5.2 Relevant hypothesis
In Table 3 and 4, we display the simulated rejection rates of Jackknife test (3.27) and the pivotal test for the hypotheses (3.25) for different values of $\Delta$ and a significance level of $5\%$. For both models the results reflect the asymptotic properties of the test described in Corollary 3.2. At the “boundary” of the hypotheses, where ${\cal M}^{2}=\Delta$, the empirical sizes of the pivotal test (4.8) are close to the nominal level $5\%$ in most cases, while the type I error of the Jackknife test (3.27) is too large for small sample sizes.
@@
In the “interior of the null hypothesis”, where $\mathcal{M}^{2}>\Delta$, the rejection rates of both tests decreases to 0 as $\Delta$ decreases and the sample size increases. Under the alternative, where $\mathcal{M}^{2}<\Delta$, the rejection rates of both tests increase to $1$ as $\Delta$ increases and the sample size increases. The rejection rates of the Jackknife test (3.27) are usually higher than those of the pivotal test (4.8).
@@
Table 4: Simulated rejection rates (in %) of Jackknife test (3.27) and the pivotal test (4.8) tests for the relevant hypotheses (3.25) in Model 2.
Summarizing, for testing for relevant deviations from sphericity from independent identically distributed data, the Jackknife test exhibits some advantages compared to the pivotal test if the sample size is sufficiently large. However, we emphasize that this observation can only be made for independent data. In the case of dependencies, the Jackknife estimator does not yield a valid testing procedure. On the other hand, a careful inspection of the proofs in the appendix shows that the asymptotic statements in Section 4 remain valid for stationary processes under appropriate mixing (Bradley,, 2007), physical dependence (Wu,, 2005) or $m$-approximability (Hörmann and Kokoszka,, 2010) conditions. Consequently, Corollary 4.1 remains valid as well, and the pivotal test (4.8) has asymptotic level $\alpha$ and is consistent for the hypotheses (3.25).
@@
Table 5: Simulated rejection rates (in %) of the Jackknife test (3.27) and the pivotal test (4.8) tests for the relevant hypotheses (3.25) in the case of dependent data.
5.3 Data analysis
For illustrate the potential applications of our approach , we analyze the log-returns of the daily exchange rate (close price) of the Yen to the Dollar and the Pound to the Euro from January 2, 2009, to December 31, 2009, which has also been investigated in Einmahl and Gantner, (2012) using a test for independent data. We downloaded the data from Yahoo Finance via the R package “quantmod” and obtained 260 entries for the log-returns, as the transactions of the currency can be conducted on weekdays including holidays. The mean of the bivariate series turns out to be $9.4\times 10^{-5}$ and $2.7\times 10^{-4}$. In Figure 1, we plot the absolute autocorrelations of different lags of both of the series. We found that there exists some evidence of serial dependence, for example, the lag 12 of the Yen to Dollar and the lag 6 of Pound to Euro.

### v2

@@ 3.2 Some statistical applications
By Proposition 3.1, Theorem 3.1 and 3.2 and Remark 3.1, the statistic $\hat{\mathcal{M}}_{n}^{2}$ is a consistent estimator of the measure $\mathcal{M}^{2}$, which defines the deviation from sphericity. To use these results for uncertainty quantification, one requires estimates of the variances $s_{n}^{2}$ and $\sigma^{2}$ defined in (3.10) and (3.12), respectively.
@@
which is obtained by replacing the integral of the squared density in (3.10) by its corresponding estimate (2.8). We will use this estimate to define a test for exact sphericity (see Remark 3.2 below).
The estimation of $\sigma^{2}$ is more difficult and we propose a Jackknife approach for this purpose (see Chapter 5 of Lee2003)). To be precise, let $\hat{\mathcal{M}}^{2}_{n-1}(-i)$ denote the estimator (2.12) of the minimum distance ${\cal M}^{2}$ based on the observations $Y_{1},\ldots,Y_{i-1},Y_{i+1},\ldots,Y_{n}$, and define the pseudovalues
@@
denotes the Jackknife estimate of the mean of $\tilde{\mathcal{M}}_{i}^{2}$.
@@
where $\Delta>0$ is a prespecified threshold. Note that this perspective of hypothesis testing aligns with the view expressed by berger1987, who argue that it is rare, and perhaps impossible, to have a null hypothesis that can be exactly modeled by a parameter being precisely 0. Similarly, tukey1991, in the context of multiple comparisons of means, emphasizes that “All we know about the world teaches us that the effects of A and B are always different — in some decimal place — for any A and B. Thus, asking ‘Are the effects different?’ is foolish.” We call hypotheses the form (3.25) and (3.26) relevant hypotheses in the following discussion. Note also that hypotheses of the form (3.26) have found considerable attention in equivalence testing in the field of biostatistics, which explains the notations $H_{0}^{\rm eq}$ and $H_{1}^{\rm eq}$ in (3.26) (see wellek2010testing) and that rejection of the null hypothesis in (3.26) allows to decide for approximate sphericity at a controlled type I error.
@@
    $\lim_{n\rightarrow\infty}\mathbb{P}\Big(\hat{\mathcal{M}}_{n}^{2}>\Delta+u_{1-\alpha}{\hat{\sigma}_{n}\over\sqrt{n}}\Big)=\begin{cases}1,&\text{if }{\mathcal{M}}^{2}>\Delta,\\ \alpha,&\text{if }{\mathcal{M}}^{2}=\Delta,\\ 0,&\text{if }{\mathcal{M}}^{2}<\Delta.\end{cases}$ (3.29)
@@
    $\lim_{n\rightarrow\infty}\mathbb{P}\Big(\hat{\mathcal{M}}_{n}^{2}\leq\Delta+u_{\alpha}{\hat{\sigma}_{n}\over\sqrt{n}}\Big)=\begin{cases}1,&\text{if }{\mathcal{M}}^{2}<\Delta,\\ \alpha,&\text{if }{\mathcal{M}}^{2}=\Delta,\\ 0,&\text{if }{\mathcal{M}}^{2}>\Delta.\end{cases}$ (3.30)
@@
    An important question in testing hypotheses of the form (3.25) and (3.26) is the choice of the threshold $\Delta$. In the following, we shall explain how to choose it in a data adaptive way for the hypotheses (3.26), and a similar argument can be given for the hypotheses (3.25). Note that these hypotheses are nested, that is, if $H_{0}^{\rm eq}$ is satisfied for $\Delta_{1}$, then it is also satisfied for all $\Delta_{2}\leq\Delta_{1}$. The decision rules for different $\Delta$’s are nested in the same way(if $H_{0}^{\rm eq}$ is rejected for $\Delta_{2}$, it is also rejected for all $\Delta_{1}\geq\Delta_{2}$). By the sequential rejection principle, we may simultaneously test the hypotheses in (3.26) for different $\Delta\geq 0$ starting at $\Delta=0$ and increasing $\Delta$ to find the minimum value of $\Delta$, say
    $\hat{\Delta}_{\alpha}:=\min\big\{\Delta\geq 0\,|\,\hat{\cal M}^{2}_{n}\leq\Delta+u_{\alpha}\hat{\sigma}_{n}/\sqrt{n}\big\}$ (3.31)
@@
    it is worthwhile mentioning that the theory developed so far also provides a test for these hypotheses without extra restrictions on the dimension $p$. More precisely, we propose to reject the null hypothesis in (3.32), whenever
@@
    where $\hat{s}_{n}$ is defined in (3.20). Then, it follows from Theorem 3.1 and 3.2 that this decision rule defines a consistent and asymptotic level $\alpha$ test. In Section 5.3, we compare the finite sample performance of the test (3.33) for the hypotheses (3.32) with other existing methods.
@@ 4 Pivotal inference
The statistical methodology developed in Section 3.2 requires the estimation of the variance $\hat{s}_{n}^{2}$. In this section, we develop a pivotal confidence interval for the deviation $\mathcal{M}^{2}$ from sphericity and pivotal tests for the hypotheses (3.25) and (3.26). For this purpose, we prove a weak convergence result for a sequential version of the estimator $\hat{\mathcal{M}}_{n}^{2}$.
@@
whose asymptotic properties are investigated in the following theorem. Throughout this section, the symbol $\Rightarrow$ denotes weak convergence in the space $\ell^{\infty}([0,1])$ of bounded functions on the interval $[0,1]$.
@@
We will now use the result (4.3) to develop pivotal inference for the measure $\mathcal{M}^{2}$. In particular, we define the statistic
@@
Similarly, we propose tests for the hypotheses (3.25) and (3.26) to reject the null hypothesis, whenever
@@ 5 Finite sample properties
In this section, we illustrate the finite sample properties of the developed methodology through a small simulation study. For the choice of kernel functions, we use the Fisher-von-Mises distribution function(2.10) for $L(\cdot)$ and Epanechnikov kernel for $K(\cdot)$, i.e., $K(x)=\frac{3}{4}(1-x^{2})$, for $|x|<1$, and $K(x)=0$ otherwise. Both of the functions satisfy the Assumption 3.1. In addition, we implement the bias correction procedure in Remark 3.1 with $a=0.9$. All results presented here are based on $1000$ simulation runs.
According to Theorem 3.2, the term $\kappa^{(p-1)/2}/(nh)$ controls the variance for the non-leading term in the Gaussian approximation, while $\sqrt{n}/\kappa^{2}+\sqrt{n}h^{4}$ accounts for the bias term with the bias reduction procedure discussed in Remark 3.2. In addition, for the sequential convergence, by Theorem 4.1, the inflated rate $(\log n)^{2}\kappa^{(p-1)/2}/(nh)$ controls the variance of the process of the non-leading term. In order to approximate Gaussianity in finite samples better, we recommend imposing a stronger control over the non-leading term, that is, $n^{3/2}\times\kappa^{(p-1)/2}/(nh)$. Therefore, we consider bandwidths $(h,\kappa)$ satisfying
@@
In the following discussion, we consider two models.
@@
Since it is difficult to obtain the analytical solution of $\mathcal{M}^{2}$, we calculate the deviation $\mathcal{M}^{2}$ numerically. We obtain $\mathcal{M}^{2}\approx 0.95$, using a larger sample size $2500$ and averaged over $20000$ times of repeated simulation. For the selection of the smoothing parameters via (5.2), we use
@@ 5.1 Confidence intervals for the deviation from a spherical distribution
For Model 1, the empirical coverage rates of both methods are close to their nominal levels for sample sizes larger than $500$. The Jackknife confidence interval achieves more accurate simulated coverage rates when the sample size is small, while the pivotal confidence intervals are more conservative for smaller sample sizes. A potential explanation for this observation is that when the sample size is small, the widths of the pivotal method are much larger than those of the Jackknife method.
@@ 5.2 Relevant hypothesis
In Table 3 and 4, we display the simulated rejection rates of Jackknife test (3.28) and the pivotal test (4.8) for the hypotheses (3.26) for different values of $\Delta$ and a significance level of $5\%$. For both models the results reflect the asymptotic properties of the test described in Corollary 3.2. At the “boundary” of the hypotheses, where ${\cal M}^{2}=\Delta$, the empirical sizes of the pivotal test (4.8) are close to the nominal level $5\%$ in most cases, while the type I error of the Jackknife test (3.28) is too large for small sample sizes.
@@
In the “interior of the null hypothesis”, where $\mathcal{M}^{2}>\Delta$, the rejection rates of both tests quickly decrease to 0 as $\Delta$ decreases and the sample size increases. Note that this is a desirable behavior for a test of the composite hypotheses (3.26) as one has to assure for a type I error that
$\sup_{{\cal M}^{2}\geq\Delta}\mathbb{P}(\text{ ``rejection''})\leq\alpha.$
Thus, the level of the test is calibrated at the “boundary” of null hypothesis ${\cal M}^{2}=\Delta$ to be close to $\alpha$. Under the alternative, where $\mathcal{M}^{2}<\Delta$, the rejection rates of both tests increase to $1$ as $\Delta$ increases and the sample size increases. The rejection rates of the Jackknife test (3.28) are usually higher than those of the pivotal test (4.8).
@@
Table 4: Simulated rejection rates (in %) of Jackknife test (3.28) and the pivotal test (4.8) for the relevant hypotheses (3.26) in Model 2.
In summary, for testing for relevant deviations from sphericity from independent identically distributed data, the Jackknife test exhibits some advantages compared to the pivotal test if the sample size is sufficiently large. However, we emphasize that this observation can only be made for independent data. In the presence of dependencies, the Jackknife estimator does not yield a valid testing procedure. On the other hand, a careful inspection of the proofs in the appendix shows that the asymptotic statements in Section 4 remain valid for stationary processes under appropriate mixing (Bradley.2007), physical dependence (Wu2005) or $m$-approximability (HrmannKok) conditions. Consequently, Corollary 4.1 remains valid as well, and the pivotal test (4.8) has asymptotic level $\alpha$ and is consistent for the hypotheses (3.26).
@@
Table 5: Simulated rejection rates (in %) of the Jackknife test (3.28) and the pivotal test (4.8) for the relevant hypotheses (3.26) in the case of dependent data.
5.3 Comparison with other methods
In this section, we compare the minimum distance approach with alternative methods for the detection of deviations from sphericity. As all other papers focus on testing for exact ellipticity or sphericity, we consider the hypotheses (3.32) and the test (3.33) in Remark 3.2. Note that under the null hypothesis ${\cal M}^{2}=0$, the estimator $\hat{\mathcal{M}}_{n}^{2}$ is unbiased (see Proposition 3.1(i)). Therefore, no bias correction as discussed in Remark 3.1 is required in this situation. For the selection of bandwidths, we balance the squared bias under the alternative hypothesis, i.e., $\kappa^{-2}+h^{4}$ and the variance under Theorem 3.1, i.e., $\kappa^{(p-1)/2}/[n(n-1)h]$, and for $p=3$ use
$h=1.5(n(n-1))^{-1/7}~,~~\kappa=1.5(n(n-1)h)^{1/3}.$ (5.7)
In the following, we compare our test with other tests for the exact hypothesis (3.32) proposed by SCHOTT2002, MPQ2002, HP2007, PG2008, babic2021 and tang2024nonparametric. We consider the null hypothesis, where $Y$ follows a $3$-dimensional standard normal distribution. For the alternative we changed the first component of $Y$ to a $(\chi_{2}^{2}-2)/2$-distribution.
In Table 6, we investigate the simulated rejection rate under the null hypothesis. The new test (3.33) provides a good approximation of the nominal level $\alpha=10\%$, $5\%$. The alternative tests exhibit a similar behavior under the null hypothesis, except for the test of tang2024nonparametric, which turns out to be conservative. The results under the alternative are displayed in Table 7. Although the new test (3.33) does not require Gaussianity or the existence of $4$-th moments, the empirical rejection rates under the alternative are larger than those of the tests HP2007, MPQ2002, SCHOTT2002, which require such assumptions, and comparable to the optimal parametric and semiparametric tests proposed by PG2008 and babic2021. Moreover, the new test (3.33) for the hypotheses (3.32) is more powerful in this scenario than the test based on kernel embedding of probabilities proposed by tang2024nonparametric.
(3.32) TL HP MPQ
[…]
200 8.6 5.0 1.0 3.7 1.7 0.3 10.2 4.4 0.5 10.9 5.5 0.5
300 10.1 5.5 1.2 3.9 1.9 0.2 10.1 5.2 1.3 10.6 5.4 1.2
400 9.2 4.6 1.4 4.1 1.5 0.3 8.0 4.0 0.5 10.0 5.9 1.1
(3.32) PG Schott BGH
[…]
200 8.6 5.0 1.0 11.0 5.1 1.2 10.1 4.8 0.6 10.9 4.7 0.9
300 10.1 5.5 1.2 10.2 5.8 1.5 9.9 4.3 0.8 11.3 5.2 1.2
400 9.2 4.6 1.4 12.2 7.2 1.3 10.8 5.4 1.2 10.9 5.9 0.9
Table 6: Simulated rejection rates (in %) of the test (3.32) and the tests of tang2024nonparametric (TL), HP2007 (HP), MPQ2002 (MPQ), PG2008 (PG), SCHOTT2002 (Schott) and babic2021 (BGH) for the null hypothesis of exact sphericity. The sample sizes are $200$, $300$ and $400$. The nominal levels are 1010%, 55% and 11%.
(3.32) TL HP MPQ
[…]
200 99.5 98.1 91.3 98.5 92.7 45.0 94.6 89.9 71.8 93.6 88.4 73.0
300 99.9 99.9 99.8 100.0 99.9 93.8 99.9 99.0 94.8 99.4 98.9 94.6
400 100.0 100.0 100.0 99.9 99.9 99.7 100.0 99.9 99.7 99.9 99.9 99.4
(3.32) PG Schott BGH
[…]
200 99.5 98.1 91.3 100.0 100.0 100.0 75.2 64.2 42.3 100.0 100.0 99.9
300 99.9 99.90 99.80 100.0 100.0 100.0 91.4 82.9 63.2 100.0 100.0 100.0
400 100.0 100.0 100.0 100.0 100.0 100.0 95.7 91.6 79.0 100.0 100.0 100.0
Table 7: Simulated rejection rates (in %) of the test (3.32) and the tests of tang2024nonparametric (TL), HP2007 (HP), MPQ2002 (MPQ), PG2008 (PG), SCHOTT2002 (Schott) and babic2021 (BGH) for the hypotheses of exact sphericity under the alternative hypothesis. The sample sizes are $200$, $300$ and $400$. The nominal levels are 1010%, 55% and 11%.
5.4 Real data example
To illustrate the potential applications of our approach, we analyze the log-returns of the daily exchange rate (close price) of the Yen to the Dollar and the Pound to the Euro from January 2, 2009, to December 31, 2009, which has also been investigated in einmahl2012 using a test for independent data. We downloaded the data from Yahoo Finance via the R package “quantmod” and obtained 260 entries for the log-returns, as the transactions of the currency can be conducted on weekdays, including holidays. The mean of the bivariate series turns out to be $9.4\times 10^{-5}$ and $2.7\times 10^{-4}$. In Figure 1, we plot the absolute autocorrelations of different lags of both of the series. We found that there exists some evidence of serial dependence, for example, the lag 12 of the Yen to Dollar and the lag 6 of the Pound to Euro.